OpenStudy (macabbey101):

HELP! WILL FAN AND MEDAL ∠DFG and ∠JKL are complementary angles. m∠DFG = x + 1, and m∠JKL = x − 9. Find the measure of each angle.

OpenStudy (anonymous):

set them equal 90 degrees because complementary means they form 90 degrees. x + 5 + x - 9 = 90 2x - 4 = 90 +4 +4 ---- ----- 2x = 90 __ __ 2 2 x = 45

OpenStudy (anonymous):

now can you message me back ;)

OpenStudy (anonymous):

how about that message